# Path Pooling: Train-Free Structure Enhancement for Efficient Knowledge Graph Retrieval-Augmented Generation

### v1

@@ Abstract
Although Large Language Models achieve strong success in many tasks, they still suffer from hallucinations and knowledge deficiencies in real-world applications. Many knowledge graph-based retrieval-augmented generation (KG-RAG) methods enhance the quality and credibility of LLMs by leveraging structure and semantic information in KGs as external knowledge bases. However, these methods struggle to effectively incorporate structure information, either incurring high computational costs or underutilizing available knowledge. Inspired by smoothing operations in graph representation learning, we propose path pooling, a simple, train-free strategy that introduces structure information through a novel path-centric pooling operation. It seamlessly integrates into existing KG-RAG methods in a plug-and-play manner, enabling richer structure information utilization. Extensive experiments demonstrate that incorporating the path pooling into the state-of-the-art KG-RAG method consistently improves performance across various settings while introducing negligible additional cost. Code is coming soon at https://github.com/hrwang00/path-pooling.
Path Pooling: Train-Free Structure Enhancement for Efficient Knowledge Graph Retrieval-Augmented Generation
@@ 1 Introduction
However, the complex knowledge structures in KGs pose challenges for effective knowledge retrieval in KG-RAG systems. Current mainstream KG-RAG methods can be categorized into path-based KG-RAG and triple-based KG-RAG based on their retrieval paradigms. Earlier path-based KG-RAG paradigm employs LLMs for path traversal on KGs Sun et al. (2024); Ma et al. (2024), subsequent path-based works try to fine-tune LLMs with KG information to retrieve knowledge graph paths Luo et al. (2024); Mavromatis and Karypis (2024). Although these methods effectively leverage the structure information of KGs, they face significant efficiency challenges, such as the high cost of path retrieval due to the vast number of possible path combinations in graph structures and the computational burden of multiple LLM calls during path traversal Sun et al. (2024); Ma et al. (2024), which limits their real-world applications. To address this efficiency issue, triple-based KG-RAG paradigm Hu et al. (2024); He et al. (2024a); Li et al. (2025) focuses on the smallest unit in KGs—the triple. Compared to the exponential number of paths in the graph, these methods significantly improve deployment efficiency. They typically use a lightweight retriever to directly retrieve query-relevant triples, followed by a single LLM call to generate responses. Notably, the latest method, SubgraphRAG Li et al. (2025), trains a simple multilayer perceptron (MLP) for triple retrieval based on semantic and structure information, achieving state-of-the-art performance. Particularly in terms of efficiency, SubgraphRAG overwhelmingly outperforms previous path-based KG-RAG methods, being over 100 times faster than the representative RoG in the retrieval process while requiring only a single LLM call for generation.
However, current triple-based KG-RAG methods, despite attempt to embed graph structure information into triple representations, primarily treat structure information as an auxiliary enhancement to semantics He et al. (2024a); Hu et al. (2024); Li et al. (2025). Many of these methods introduce complex structure embedding models, which significantly increase training complexity and cost. At the same time, due to the inherent nature of triples as the atomic units of a knowledge graph, they inevitably fall short in capturing structure information compared to path-based KG-RAG methods, where paths inherently encode richer structure context. This raises a critical question: Can we leverage graph structure for KG-RAG in a low-cost, train-free manner within the triple-based paradigm?
@@ 3.2 Triple-based KG-RAG
Triple-based KG-RAG methods retrieve a triple sequence $T$ from KGs by leveraging semantic and structure information of KGs jointly. In terms of semantic information, triple-based KG-RAG methods initiate the retrieval process by encoding both the query and information of KGs into a unified vector space. They typically apply an off-the-shelf text encoder (e.g., SentenceBert Reimers and Gurevych (2019); gte-large-en-v1.5 Zhang et al. (2024c)) to embed the textual attributes of entities and relations within the KGs, yielding the corresponding embeddings. The same encoding strategy is employed to the query in order to ensure consistency. Next, to extract the most relevant entities and relations in the KGs, triple-based KG-RAG methods employ a scorer (e.g., cosine similarity; MLP) to match semantic similarity between them and the given query. In terms of structure information, GNNs and its variant are widely used to effective capture topology structure of KGs. SubgraphRAG utilizes structure information as feature Li et al. (2025) to facilitate the aforementioned semantic matching process, while other methods He et al. (2024a); Hu et al. (2024) align structure information with extracted text information to augment the LLMs. The above retrieval process essentially functions as a pruning operation, preserving a compact triple sequence $T$ most relevant to given query for subsequently generation process. $T$ can complement the missing knowledge of LLMs. Algorithm 2 describes the workflow of triple-based KG-RAG, where path pooling effectively enhance graph structure information in a plug-and-play manner.
@@ 4.1 Experiment Setup
Implementation Details. We enhance the state-of-the-art triple-based KG-RAG method, SubgraphRAG, using path pooling through positional reranking and triple reselection. For brevity, we abbreviate SubgraphRAG as S-RAG in the following experiments. To evaluate the performance of our path pooling strategy, we employ a diverse selection of open-source LLMs: Llama 3.1-8B, Llama 3.1-70B Llama Team (2024), Qwen 2.5-7B, and Qwen 2.5-72B Qwen et al. (2025). For all models, we use a one-shot prompt for reasoning, as detailed in Appendix A.4. To ensure reproducibility, we set the temperature to 0 and the maximum token length for generation to 4000, aligning with previous research Li et al. (2025).
@@ 4.2 Main Results
The experimental results of the reranking and reselection mechanism, which refine the triple sequence using path pooling, are presented in Table 3. Overall, reranking introduces minimal additional retrieval time, as it only requires reordering triples based on the smoothed scores from path pooling. For instance, even when processing up to 200 triples, the additional retrieval time is only 3.2 ms per query, while achieving a 3.1% improvement in Hit@1 and a 2.3% increase in F1-Score on Qwen2.5-7b. Furthermore, when applying path pooling for triple reselection, the performance gains become even more pronounced. We conduct experiments by first retrieving 500 triples using SubgraphRAG and then selecting 25 to 200 triples based on their path pooling scores. Notably, with just 25 triples, the reselection mechanism boosts Hit@1 by 5.2% on Llama3.1-70b and 5.3% on Qwen2.5-7b while introducing only 18.4ms per query. Remarkably, using only 25 triples, the reselection mechanism even outperforms the original SubgraphRAG utilizing 200 triples, demonstrating its high effectiveness.
Figure 2 provides a detailed visualization of this phenomenon in triple number 50. Our path pooling strategy consistently enhances the performance of SubgraphRAG across various settings. Simply leveraging the structure information from path pooling to rerank triples in the prompt based on positional bias already leads to notable improvements. Moreover, refining triple reselection using path pooling scores to filter out irrelevant noise results in further performance gains. Figure 3 further visualizes the performance of Llama 3.1-8B  [^3] In the paper of SubgraphRAG, the Hit of Llama3.1-8b with 100 triples is 56.98. as the number of triples increases from 25 to 200. When using path pooling for reranking, performance consistently improves with more triples, achieving substantial gains at 200 triples, with Hit@1 and Macro-F1 increasing by. This aligns with prior findings on positional bias: the longer the sequence, the more pronounced the bias Liu et al. (2023). In contrast, when applying triple reselection, path pooling achieves significant improvements with as few as 25 triples, boosting Hit@1 and Macro-F1 by. However, as the number of triples increases further, the performance gains gradually diminish.
@@ 4.3.1 Different Path Search Algorithms
Beyond the widely used shortest path algorithm, Dijkstra, other graph traversal methods such as BFS and Random Walk can also be employed to extract structure information. We evaluate the impact of different path search algorithms within our path pooling strategy on Llama3.1-8b, as shown in Figure 4. To prevent excessive smoothing of structure information, we set the maximum path length for BFS and Random Walk to 4 , and Random Walk samples 256 times. Overall, all path search algorithms contribute to performance improvements when integrated into path pooling. However, Random Walk performs significantly worse than the other two, likely because its stochastic nature fails to capture key structure information. While BFS and Dijkstra achieve similar performance, BFS introduces a significantly larger number of paths, leading to increased computational overhead, as shown in Table 2. For instance, with 500 triples, BFS requires 232.2 ms, whereas Dijkstra only takes 18.1 ms, making it 12.8x times more efficient. Thus, we recommend using Dijkstra in the path pooling strategy to achieve a balance between effectiveness and efficiency.
@@ 4.3.2 Different Pooling Strategies
In Figure 5, we further compare the performance of different smoothing operations in our path pooling strategy. We evaluate two commonly used pooling operation for smoothing: max pooling, which retains the maximum score along the path, and average pooling, which computes the mean score of triples on the path as the base score. Both operation could capture local graph structure information, enhancing the reasoning performance of LLMs. However, average pooling consistently outperforms max pooling, especially when the number of triples is small. For instance, with 50 triples, average pooling shows a significant advantage over max pooling. Using max pooling provides a increase of 1.27%, from 50.86 to 49.59, while leveraging average pooling brings a substantial improvement of 4.42%, from 54.01 to 49.59. This suggests that average pooling better preserves neighborhood information, making it the preferred choice for smoothing in path pooling.
@@ 4.3.3 Lost in the Middle Positional Bias
Figure 6: Reranking Results of Lost in the Middle on Qwen2.5-7b and Llama3.1-8b
@@ 5 Related Work
Triple-based KG-RAG retrieves the top-$k$ triples relevant to the query, which is more efficient compared to path-based KG-RAG. GRAG Hu et al. (2024) employs a divide-and-conquer strategy to retrieve subgraphs that are most relevant to the query.During the reasoning phase, GRAG provides the LLMs with both text and graph views of information, enabling a deeper understanding of the relations between entities. G-retrieval He et al. (2024a) integrates the advantages of LLM, GNN and RAG, employing cosine similarity to retrieve entities and relations, while the PCST algorithm manages the size of the subgraph. The textual representations of the subgraph, along with the structure information outputted by the GNN, are fed into the LLM to generate answers. Unlike previous work, currently SubgraphRAG Li et al. (2025) trains a lightweight multilayer perceptron, which enables scalable and flexible retrieval process, achieving higher retrieval efficiency and better performance. Compared to path-based KG-RAG, none of the above methods require fine-tuning or multiple calls to the LLMs. Our path pooling strategy can be seamlessly pluged into the retrieval and generation phases of the triple-based KG-RAG in a low-cost, train-free manner. It enhances the performance of LLMs by introducing graph structure information in the retrieved triple sequence.
@@ 6 Conclusion
In this paper, we address a key challenge in KG-RAG: integrating rich structure information from knowledge graphs without incurring excessive computational costs. Current triple-based KG-RAG methods are efficient but fail to capture the full graph structure, while path-based KG-RAG methods, though more expressive, are computationally expensive. To bridge this gap, we propose the path pooling strategy to enhance existing triple-based KG-RAG in a plug-and-play manner by leveraging graph smoothing effects. Our method identifies key paths using graph search and applies smoothing to refine triple scores along these paths, effectively enriching the structure representation in triple-based KG-RAG. Experiments demonstrate that path pooling consistently improves KG-RAG performance with minimal overhead. This work introduces a novel train-free enhancement strategy for KG-RAG by leveraging smoothing effects widely used in graph learning. Future work could explore more effective path searching and finer-grained smoothing techniques to capture richer structure information.
@@ A.1.2 Different Pooling Strategies
On Qwen2.5-7b, triple reselection mechanism with different pooling strategies still improves SubgraphRAG, as shown in Figure 8. Notably, SubgraphRAG with max pooling can achieve better results than the original SubgraphRAG across all triple numbers, both Hit@1 and F1-score. Max pooling can provide a smoothing effect for triples along the path. However, since it retains only the maximum along that path, it may loss some important neighborhood information. In contrast, average pooling demonstrates better improvement compared to max pooling across various parameter settings. For example, max pooling provides a 1.84% improvement over not using path pooling from 45.23 to 47.07, while average pooling brings 3.06% more improvement than max pooling from 47.07 to 50.13 with 50 triples on Hit@1, demonstrating its significant effectiveness.
@@ A.2.2 Different Pooling Strategies
Next, we present the experimental results of the positional reranking mechanism with path pooling.
Triple Num 500   +18.1ms/query
Methods F1-Score Hit@1
w/o ours w/ ours $\Delta$ w/o ours w/ ours $\Delta$
Llama-8b 45.5 48.3 +2.8 49.2 53.3 +4.1
Llama-70b 51.8 53.6 +1.8 56.9 58.7 +1.8
Qwen-7b 41.7 43.5 +1.8 44.9 47.1 +2.2
Qwen-72b 52.2 54.7 +2.5 55.8 58.8 +3.0
Table 3: Reranking Results with 500 triples
When the number of triples reaches 500, the increase in token counts results in a long sequence reasoning problem. However, SubgraphRAG enhanced with our positional reranking mechanism still demonstrates substantial advantages. For instance, this is particularly evident in the performance of Llama3.1-8b, as shown in Table 3. Leveraging our positional reranking mechanism achieves a 4.1% improvement in Hit@1 and a 2.8% increase in F1-Score.

### v2

# Path Pooling: Training-Free Structure Enhancement for Efficient Knowledge Graph Retrieval-Augmented Generation

@@ Abstract
Although Large Language Models achieve strong success in many tasks, they still suffer from hallucinations and knowledge deficiencies in real-world applications. Many knowledge graph-based retrieval-augmented generation (KG-RAG) methods enhance the quality and credibility of LLMs by leveraging structure and semantic information in KGs as external knowledge bases. However, these methods struggle to effectively incorporate structure information, either incurring high computational costs or underutilizing available knowledge. Inspired by smoothing operations in graph representation learning, we propose path pooling, a simple, train-free strategy that introduces structure information through a novel path-centric pooling operation. It seamlessly integrates into existing KG-RAG methods in a plug-and-play manner, enabling richer structure information utilization. Extensive experiments demonstrate that incorporating the path pooling into the state-of-the-art KG-RAG method consistently improves performance across various settings while introducing negligible additional cost.
Path Pooling: Training-Free Structure Enhancement for Efficient Knowledge Graph Retrieval-Augmented Generation
@@ 1 Introduction
However, the complex knowledge structures in KGs pose challenges for effective knowledge retrieval in KG-RAG systems. Current mainstream KG-RAG methods can be categorized into path-based KG-RAG and triple-based KG-RAG based on their retrieval paradigms. Earlier path-based KG-RAG paradigm employs LLMs for path traversal on KGs Sun et al. (2024); Ma et al. (2024), subsequent path-based works try to fine-tune LLMs with KG information to retrieval knowledge graph paths Luo et al. (2024); Mavromatis and Karypis (2024). Although these methods effectively leverage the structure information of KGs, they face significant efficiency challenges, such as the high cost of path retrieval due to the vast number of possible path combinations in graph structures and the computational burden of multiple LLM calls during path traversal Sun et al. (2024); Ma et al. (2024), which limits their real-world applications. To address this efficiency issue, triple-based KG-RAG paradigm Hu et al. (2024); He et al. (2024a); Li et al. (2025) focuses on the smallest unit in KGs—the triple. Compared to the exponential number of paths in the graph, these methods significantly improve deployment efficiency. They typically use a lightweight retriever to directly retrieve query-relevant triples, followed by a single LLM call to generate responses. Notably, the latest method, SubgraphRAG Li et al. (2025), trains a simple multilayer perceptron (MLP) for triple retrieval based on semantic and structure information, achieving state-of-the-art performance. Particularly in terms of efficiency, SubgraphRAG overwhelmingly outperforms previous path-based KG-RAG methods, being over 100 times faster than the representative RoG in the retrieval process while requiring only a single LLM call for generation.
However, current triple-based KG-RAG methods, despite attempt to embed graph structure information into triple representations, primarily treat structure information as an auxiliary enhancement to semantics He et al. (2024a); Hu et al. (2024); Li et al. (2025). Many of these methods introduce complex structure embedding models, which significantly increase training complexity and cost. At the same time, due to the inherent nature of triples as the atomic units of a knowledge graph, they inevitably fall short in capturing structure information compared to path-based methods, where paths inherently encode richer structure context. This raises a critical question: Can we leverage graph structure for KG-RAG in a low-cost, training-free manner within the triple-based paradigm?
@@ 3.2 Triple-based KG-RAG
Triple-based KG-RAG methods retrieve a triple sequence $T$ from KGs by leveraging semantic and structure information of KGs jointly. In terms of semantic information, triple-based KG-RAG methods initiate the retrieval process by encoding both the query and information of KGs into a unified vector space. They typically apply an off-the-shelf text encoder (e.g., SentenceBert Reimers and Gurevych (2019); gte-large-en-v1.5 Zhang et al. (2024c)) to embed the textual attributes of entities and relations within the KGs, yielding the corresponding embeddings. The same encoding strategy is employed to the query in order to ensure consistency. Next, to extract the most relevant entities and relations in the KGs, triple-based KG-RAG methods employ a scorer (e.g., cosine similarity; MLP) to match semantic similarity between them and the given query. In terms of structure information, GNNs and its variant are widely used to effective capture topology structure of KGs. SubgraphRAG utilizes structure information as feature to facilitate the aforementioned semantic matching process, while other methods align structure information with extracted text information to augment the LLMs. The above retrieval process essentially functions as a pruning operation, preserving a compact triple sequence $T$ most relevant to given query for subsequently generation process. $T$ can complement the missing knowledge of LLMs. Algorithm 2 describes the workflow of triple-based KG-RAG, where path pooling effectively enhance graph structure information in a plug-and-play manner.
@@ 4.1 Experiment Setup
Implementation Details. We enhance the state-of-the-art triple-based KG-RAG method, SubgraphRAG, using path pooling through positional reranking and triple reselection. For brevity, we abbreviate SubgraphRAG as S-RAG in the following experiments. To evaluate the performance of our path pooling strategy, we employ a diverse selection of open-source LLMs: Llama3.1-8b-Instruct, Llama3.1-70b-Instruct Llama Team (2024), Qwen2.5-7b-Instruct, and Qwen2.5-72b-Instruct Qwen et al. (2025). For all models, we use a one-shot prompt for reasoning, as detailed in Appendix A.4. For all LLMs, we employed one-shot prompt for reasoning. The reasoning prompt is detailed in Appendix A.4. To ensure reproducibility, we set the temperature to 0 and the maximum token length for generation to 4000, aligning with previous research Li et al. (2025).
@@ 4.2 Main Results
The experimental results of the reranking mechanism, which refines the triple sequence using path pooling, are presented in Table 3. Overall, reranking introduces minimal additional retrieval time, as it only requires reordering triples based on the refined scores from path pooling. For instance, even when processing up to 200 triples, the additional retrieval time is only 3.2 ms per query, while achieving a 3.1% improvement in Hit@1 and a 2.3% increase in F1-Score on Qwen2.5-7b. Furthermore, when applying path pooling for triple reselection, the performance gains become even more pronounced. We conduct experiments by first retrieving 500 triples using SubgraphRAG and then selecting 25 to 200 triples based on their path pooling scores. Notably, with just 25 triples, the reselection mechanism boosts Hit@1 by 5.2% on Llama3.1-70b and 5.3% on Qwen2.5-7b while introducing only 18.4ms per query. Remarkably, using only 25 triples, the reselection mechanism even outperforms the original SubgraphRAG utilizing 200 triples, demonstrating its high effectiveness.
Figure 2 provides a detailed visualization of this phenomenon in triple num 50. Our path pooling strategy consistently enhances the performance of SubgraphRAG across various settings. Simply leveraging the structure information from path pooling to rerank triples in the prompt based on positional bias already leads to notable improvements. Moreover, refining triple reselection using path pooling scores to filter out irrelevant noise results in further performance gains. Figure 3 further visualizes the performance of Llama 3.1-8B  [^3] In the paper of SubgraphRAG, the Hit@1 of Llama3.1-8b-Instruct with 100 triples is 56.98. as the number of triples increases from 25 to 200. When using path pooling for reranking, performance consistently improves with more triples, achieving substantial gains at 200 triples, with Hit@1 and Macro-F1 increasing by. This aligns with prior findings on positional bias: the longer the sequence, the more pronounced the bias Liu et al. (2023). In contrast, when applying triple reselection, path pooling achieves significant improvements with as few as 25 triples, boosting Hit@1 and Macro-F1 by. However, as the number of triples increases further, the performance gains gradually diminish.
@@ 4.3.1 Different Path Search Algorithms
Beyond the widely used shortest path algorithm, Dijkstra, other graph traversal methods such as BFS and Random Walk can also be employed to extract structure information. We evaluate the impact of different path search algorithms within our path pooling strategy on Llama3.1-8b, as shown in Figure 4. To prevent excessive smoothing of structural information, we set the maximum path length for BFS and Random Walk to 4, and Random Walk samples 256 times. Overall, all path search algorithms contribute to performance improvements when integrated into path pooling. However, Random Walk performs significantly worse than the other two, likely because its stochastic nature fails to capture key structure information. While BFS and Dijkstra achieve similar performance, BFS introduces a significantly larger number of paths, leading to increased computational overhead, as shown in Table 2. For instance, with 500 triples, BFS requires 232.2 ms, whereas Dijkstra only takes 18.1 ms, making it 12.8x times more efficient. Thus, we recommend using Dijkstra in the path pooling strategy to achieve a balance between effectiveness and efficiency.
@@ 4.3.2 Different Pooling Strategies
In Figure 5, we further compare the performance of different smoothing operations in our path pooling strategy. We evaluate two commonly used pooling operation for smoothing: max pooling, which retains the maximum score along the path, and average pooling, which computes the mean score of triples on the path as the base score. Both operation could capture local graph structure information, enhancing the reasoning performance of LLMs. However, average pooling consistently outperforms max pooling, especially when the number of triples is small. For instance, with 50 triples, average pooling shows a significant advantage over max pooling. Using max pooling provides a increase of 1.27%, from 0.5086 to 0.4959, while leveraging average pooling brings a substantial improvement of 4.42%, from 0.5401 to 0.4959. This suggests that average pooling better preserves neighborhood information, making it the preferred choice for smoothing in path pooling.
@@ 4.3.3 Lost in the Middle Positional Bias
Figure 6: Reranking Results of Lost in Middle on Llama3.1-8b
@@
4.3.4 Positional Reranking with More Triples
Triple Num 500   +18.1ms/query
Methods F1-Score Hit@1
w/o ours w/ PR $\Delta$ w/o ours w/ PR $\Delta$
Llama-8b 45.5 48.3 +2.8 49.2 53.3 +4.1
Llama-70b 51.8 53.6 +1.8 56.9 58.7 +1.8
Qwen-7b 41.7 43.5 +1.8 44.9 47.1 +2.2
Qwen-72b 52.2 54.7 +2.5 55.8 58.8 +3.0
Table 3: Reranking Results with 500 triples
Table 3 presents the results of the positional reranking mechanism with path pooling using more triples. As the number of triples expands, for instance to 500, the corresponding increase in token counts introduces challenges associated with long sequence reasoning. Despite this, SubgraphRAG, when augmented with our positional reranking mechanism, demonstrates notable performance advantages. This is particularly evident in the results for Llama3.1-8b. Specifically, leveraging our positional reranking mechanism yields a 4.1% improvement in Hit@1 and a 2.8% increase in F1-Score.
@@ 5 Related Work
Triple-based KG-RAG retrieves the top-$k$ triples relevant to the query, which is more efficient compared to path-based KG-RAG. GRAG Hu et al. (2024) employs a divide-and-conquer strategy to retrieve subgraphs that are most relevant to the query, utilizing a graph pruning mechanism to unify and refine them. During the reasoning phase, GRAG provides the LLMs with both text and graph views of information, enabling a deeper understanding of the relations between entities. G-retrieval He et al. (2024a) integrates the advantages of LLM, GNN and RAG, which employs cosine similarity to retrieve entities and relations, while the PCST algorithm manages the size of the subgraph. The textual representations of the subgraph, along with the structure information outputted by the GNN, are fed into the LLM to generate answers. Unlike previous work, currently SubgraphRAG Li et al. (2025) trains a lightweight multilayer perceptron, which enables scalable and flexible retrieval processes, achieving higher retrieval efficiency and better performance. Compared to path-based KG-RAG, none of the above methods require fine-tuning or multiple calls to the LLMs. Our path pooling strategy can be seamlessly pluged into the retrieval and generation phases of the triple-based KG-RAG in a low-cost, training-free manner. It enhances the performance of LLMs by introducing graph structure information in the retrieved triple sequence.
@@ 6 Conclusion
In this paper, we address a key challenge in KG-RAG: integrating rich structure information from knowledge graphs without incurring excessive computational costs. Current triple-based KG-RAG methods are efficient but fail to capture the full graph structure, while path-based KG-RAG methods, though more expressive, are computationally expensive. To bridge this gap, we propose the path pooling strategy to enhance existing triple-based KG-RAG in a plug-and-play manner by leveraging graph smoothing effects. Our method identifies key paths using graph search (e.g., Dijkstra’s algorithm) and applies smoothing to refine triple scores along these paths, effectively enriching the structure representation in triple-based KG-RAG. Experiments demonstrate that path pooling consistently improves KG-RAG performance with minimal overhead. This work introduces a novel training-free enhancement strategy for KG-RAG by leveraging smoothing effects widely used in graph learning. Future work could explore more effective path searching and finer-grained smoothing techniques to capture richer structure information.
Limitations
While path pooling effectively leverages paths searched from query entities to smooth retrieved triples, thereby addressing the underutilization of structural information in KG-RAG, its full potential remains underexplored. Our current experiments predominantly relied on conventional path search algorithms, such as BFS or Dijsktra, without investigating potentially more efficacious weighted or heuristic path searching methods. Furthermore, introducing structural information via paths for triple smoothing represents only one specific method. KGs possess richer structural characteristics, such as node degree, adjacency relationships, and local subgraph topologies, which offer promising yet unexplored avenues for effective triple smoothing and warrant future investigation.
@@ A.1.2 Different Pooling Strategies
On Qwen2.5-7b, triple reselection mechanism with different pooling strategies still improves SubgraphRAG, as shown in Figure 8. Notably, SubgraphRAG with max pooling can achieve better results than the original SubgraphRAG across all triple numbers, both Hit@1 and F1-score. Max pooling can provide a smoothing effect for triples along the path. However, since it retains only the maximum along that path, it may loss some important neighborhood information. In contrast, average pooling demonstrates better improvement compared to max pooling across various parameter settings. For example, max pooling provides a 1.84% improvement over not using path pooling from 45.23 to 47.07, while average pooling brings 3.06% more improvement than max pooling from 47.07 to 50.13 with 50 triples, demonstrating its significant effectiveness.